OpenStudy (anonymous):

John wants to rent a tent to go camping. The cost to rent a tent from Woodland Outfitters is a flat rate of $10.00 plus $0.50 for each night the tent is rented. Part a. Let x represent the number of nights John keeps the tent and write a cost function, c(x), for the cost to rent a tent from Woodland Outfitters. Part b. How much will John owe if he uses the tent for five nights?

OpenStudy (anonymous):

first You should write your own function. If you have to pay $10 just to get a tent without even using it (you used it for 0 nights), then the y-intercept is 10. For each night, you pay $0.50 more dollars (or 1/2). This is the slope. The function is c(x) = (1/2)x + 10.

OpenStudy (anonymous):

Then, if you plug in 5 for x, c(5) = (1/2)(5) + 10 c(5) = (5/2) + 10 c(5) = $12.50

OpenStudy (anonymous):

thats part a :) thank you. and part b will be 0.50(5) + 10 = 12.50

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

thank you :)